OpenStudy (anonymous):

write the slope intercept equation line pass through -12,10 but perpendicular to 4x+6y=3

OpenStudy (anonymous):

okay so first find gradient/slope of this line 6y=3-4x y=3/6-4x/6 slope=-4/6 A line perpendicular to this will have -1/slope of the line it is perpendicular to slope of perpendicular=6/4=3/2 it passes through points -12,10 using equation formula y-10=3/2(x+12) 2y-20=3x+36 answer